OpenStudy (mendicant_bias):

Could somebody explain linear independence to me? I was asked earlier to show whether not the solutions to some (arbitrary) DE are x, x+1, and x^2, and I wasn't sure what to do or how to approach it.

OpenStudy (perl):

would need more information to help you

OpenStudy (zzr0ck3r):

A set of elements of some vector space are linearly independent if no one of the elements can be written as a linear combination of the other elements. example: Consider the subset of \(\mathbb{R}^3\) \(\{(1,2,3),(4,5,6),(9,12,15)\} \). Lets name these vectors \(a,b\) and \(c\) respectively. This set is NOT linearly independent because \(a+2b=c\). We might as well remove \(c\), because we can "get to it" through the other two vectors. then \(\{a,b\}\) is linearly independent. because \(1*4 = 4\) but \(2*4 \ne 5\). hope this helps.

OpenStudy (mendicant_bias):

So, none of these are linearly *dependent*, because any constant multiple will not achieve one equalling the other, correct?

OpenStudy (mendicant_bias):

e.g. they are all linearly dependent?

OpenStudy (mendicant_bias):

(The x, x+1, and x^2)

OpenStudy (zzr0ck3r):

so you are in R[x]?

OpenStudy (mendicant_bias):

I believe so, but don't really know, I haven't taken Linear Algebra-which sounds relevant-and I'm dealing with sets of solutions to ODE's, e.g. all three of those are proposed as solutions, but whether they are linearly independent or not is what's in question. I don't know if them being ODE's affects the idea of linear independence or not.

OpenStudy (zzr0ck3r):

wih polys of max degree 2?

OpenStudy (zzr0ck3r):

they are, you cant make any one of these polynomials by adding any of the other polynomials together and multiplying by constants.

OpenStudy (mendicant_bias):

Alright; how does the polynomial degree affect the situation?

OpenStudy (zzr0ck3r):

nothing really, I just wanted to make sure we were talking about the same thing...

OpenStudy (mendicant_bias):

Lol, yeah, we are. Thanks so much.